Stock J has a beta of 1.23 and an expected return of 13.25 percent, while Stock K has a beta of .84 and an expected return of 10.60 percent. You want a portfolio with the same risk as the market. a. What is the portfolio weight of each stock? (Do not round intermediate calculations and round your answers to 4 decimal places, e.g., .1616.) b. What is the expected return of your portfolio? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

a. Beta of Market =1
Let Weight of Stock j =x
Weight of Stock K =1-x
Weight of Stock J*Beta of J+Weight of Stock K*Beta of K =1
x*1.23+(1-x)*0.84 =1
x*(1.23-0.84)=1-0.84
Portfolio Weight of J x =0.16/(1.23-0.84) =41.03%
Portfolio Weight of K ,(1-x) =1-41.03% =58.87%

b. Expected Return of Portfolio =13.25%*41.03%+10.60%*58.87% =11.69%